# LOCUS – LOcalization with Channel Uncertainty and Sporadic Energy

### v2

@@ Abstract.
Accurate sound source localization (SSL) requires consistent multichannel data for reliable degree of arrival (DoA) estimation. However, intermittently powered batteryless systems often suffer from incomplete sensor data due to the stochastic nature of energy harvesting. Existing methods struggle with missing channels, leading to significant performance degradation. In this paper, we propose LOCUS, a novel deep learning-based system designed to recover corrupted features for SSL in batteryless systems. LOCUS addresses missing data by leveraging information entropy estimation and conditional interpolation, combining three modules: (1) Information-Weighted Focus (InFo), which identifies and quantifies corrupted data elements, (2) Latent Feature Synthesizer (LaFS), which synthesizes missing features, and (3) Guided Replacement (GRep), which intelligently replaces missing elements while preserving valid data. We demonstrate significant performance improvements using two datasets: DCASE and LargeSet, where LOCUS achieves up to 36.91% lower DoA error compared to existing methods. Real-world evaluations across three environments with intermittent power sources show a 25.87-59.46% improvement in performance when channels are stochastically missing. Additionally, we release a 50-hour multichannel dataset to support further research in SSL.
@@ 1. Introduction
Accurate degree of arrival (DoA) estimation relies on extracting temporal and spatial differences from multichannel data captured by multiple microphones. Both signal processing (Schmidt, 1986; DiBiase, 2000; Roy and Kailath, 1989) and deep learning (Xiao et al., 2015; Yalta et al., 2017; Chakrabarty and Habets, 2017) approaches use this multichannel data to localize sound sources effectively. In real-world scenarios, where complex multipath profiles and room impulse responses arise from signal reflections and refractions, deep learning methods excel by leveraging the richness of multichannel data to capture intricate patterns that traditional signal processing techniques may miss.
However, intermittently powered batteryless systems face significant challenges in maintaining full-channel data integrity due to the stochastic nature of energy harvesting. These systems, reliant on sporadic power availability, often struggle to consistently capture data from all channels, which is crucial for precise DoA estimation. Intermittent systems typically can only process short-lived events, lasting only a few hundred milliseconds, and are primarily limited to tasks such as classification (Majid et al., 2020a; Islam and Nirjon, 2020; Monjur et al., 2023). This constraint arises from the core design of intermittent systems, which rely on a fixed-size capacitor to store sufficient harvested energy for operation. During the system’s active phase, it consumes this stored energy to perform designated tasks, after which it enters a sleep phase when the energy is depleted. By selecting different capacitor sizes, the duration of the active and sleep phases can be adjusted. However, intermittent systems inevitably encounter one or more sleep phases when insufficient energy is harvested, resulting in incomplete or missing sensor data (Monjur et al., 2023). This lack of sensor data significantly degrades the system’s performance, particularly for longer events that last on the order of tens of seconds. As a result, these systems are not typically used for sound source localization (SSL), where complete multichannel data is essential.
Despite these challenges, batteryless systems hold great potential in applications such as environmental monitoring and wildlife tracking, where low-power, autonomous operation is essential. Even though these systems experience frequent power interruptions, they can still provide valuable data during operational periods.
Working with features instead of raw time-domain data is particularly important in batteryless systems. Extracted features like generalized cross-correlation (GCC) and spectral power encapsulate the most relevant information for sound source localization, reducing noise and data dimensionality. This not only improves computational efficiency but also minimizes energy consumption, which is critical in resource-constrained systems. Moreover, feature extraction can be implemented using low-power analog circuits in embedded systems, further reducing the energy footprint.
Existing solutions for addressing missing or corrupted data in sound source localization typically rely on statistical imputation techniques such as mean imputation (Allison, 2001), hot deck imputation (Little and Rubin, 2019), and multiple imputations (Rubin, 1996). While these methods can fill data gaps, they have notable limitations. Mean imputation oversimplifies by averaging values, hot deck imputation can introduce inaccuracies if data is poorly matched, and multiple imputations are computationally intensive and struggle with long missing data sequences (Hasan et al., 2021). Recent approaches using mutual information and conditional entropy (Vergara and Estévez, 2014) help identify corrupted data but fall short in recovering the information necessary for DNN-based sound source localization. Additionally, recent work on imputation-based data recovery integrated into DNN pipelines for batteryless systems (Monjur et al., 2023) is designed for single-channel input, making it inadequate for the multi-channel sound source localization needed in more complex systems.
The gap in current sound source localization systems lies in the inability to effectively recover missing channel data caused by power interruptions in intermittently powered systems. Without advanced methods to intelligently infer or reconstruct this information, localization accuracy suffers significantly. Frequent data loss, especially in batteryless systems relying on stochastic energy harvesting, presents a major challenge for precise localization.
This paper bridges the gap between faulty multi-channel microphone arrays and deep learning-based batteryless sound source localization (SSL) by proposing LOCUS. LOCUS combines information entropy estimation and conditional interpolation to identify corrupted portions in the feature set and estimate the true feature values. The system consists of three novel modules: (1) information-weighted focus (InFo), which estimates the information entropy for each element of the calculated features to understand the impact of corrupted data; (2) latent feature synthesizer (LaFS), which estimates missing information by leveraging the relationships derived from information entropy; and (3) guided replacement (GRep), which uses the information entropy from InFo to replace corrupted elements with interpolated features from LaFS, leaving unaffected elements intact. These reconstructed features are then fed into the DNN for sound source localization.
The main contributions of this work are as follows:
  • We propose LOCUS, a three-step information recovery system that leverages information entropy estimation and conditional interpolation to recover corrupted features, mitigating performance loss in deep learning-based sound source localization due to missing or faulty input channels.
  • We provide a mathematical formulation of the problem and establish a connection with information theory principles.
  • Beyond evaluation with multiple datasets, we conduct real-world evaluations in three different environments, where missing channels occur due to power intermittence from energy harvesting systems (solar and radio frequency). This demonstrates the practical applicability and robustness of LOCUS in real-world scenarios.
  • We release a new 6-channel acoustic dataset (LargeSet) comprising 50 hours of audio recorded in 10 different environments. This dataset supports further research and development in sound source localization.
Our evaluation shows significant performance gains, with LOCUS achieving 21.27% and 17.14% less average DoA error than $SoundSeive$(Monjur et al., 2023) on the DCASE (Politis et al., 2021) and LargeSet datasets, respectively. In real-world evaluations, LOCUS improves performance by 25.87-59.46% in scenarios with stochastically missing information. Additionally, LOCUS reduces the 90th percentile localization error by 25.52% in wireless source localization tasks using the WiFi channel state information (CSI) dataset (Ayyalasomayajula et al., 2020).
2. Managing Missing Data in Batteryless Microphone Arrays
Quantifying Missing Data (MDP). In a multi-channel microphone array, data from a sound source can be corrupted by missing or intermittent data from one or more channels, while others may remain unaffected. To standardize the measurement of missing data, we define the Missing Data Percentage (MDP), denoted as $p\%$. The MDP quantifies the proportion of the total audio that is corrupted. The MDP is calculated using the following formula:
@@ 2. Managing Missing Data in Batteryless Microphone Arrays
where $t_{1},t_{2},\ldots,t_{n}$ represent the segments with missing data, and $T$ is the total audio duration.
For example, Figure 1 illustrates an instance with a 75% MDP, where up to two channels are missing at any given time. In time segments $t_{1}$ and $t_{2}$, which account for 75% of the total duration, one or more channels are missing data. In contrast, segment $t_{3}$ has no missing data across any channels. As a result, the MDP in this example is 75%.
Figure 1. Example of raw audio with 75% missing data.
Figure 2. Example of the missing channels with RF harvester.
Missing Channels in Intermittent Systems. Intermittent power systems, such as those powered by RF energy harvesters, frequently experience unpredictable power fluctuations that can lead to significant data loss in multi-channel microphone arrays. These fluctuations are caused by external factors, such as uncontrolled human traffic between the RF transmitter and receiver, which dynamically affects the system’s ability to maintain consistent power. As a result, some microphones intermittently fail to collect data, leading to incomplete or corrupted information across the array.
Figure 2 illustrates energy traces from an RF-harvested system over a 120-second period, showing how fluctuations in available energy result in a 56.67% MDP. This scenario demonstrates how missing data affects nearly 60% of the total audio, with varying numbers of channels missing at any given time.
In Section 7, we evaluate the impact of stochastic power variations on sound source localization. The results highlight the need for robust techniques like LOCUS to handle missing data in real-world applications, where power instability is common, and traditional methods often fail.
Impact of Missing Channels on SSL. There is extensive research on sound source localization (SSL) with multichannel microphones, both in simulated and real-world environments (Evers et al., 2020; Tan et al., 2021). However, while some studies address noise and reverberation, the impact of missing channel data remains underexplored.
Popular SSL input features such as Mel-frequency cepstral coefficients (MFCC) and Generalized Cross-Correlation (GCC) are highly sensitive to missing channel data, distorting feature distributions and degrading DNN performance. Figure 3 shows how MFCC and GCC change with and without missing data. It is evident how missing information distorts the features and makes them less reliable for downstream DNN tasks.To quantify this effect, we evaluate an existing SSL algorithm (Shimada et al., 2021; Adavanne et al., 2018) on the two datasets (Section 5.1), with varying levels of MDP. Figure 4 shows how DoA error increases as MDP rises, emphasizing the need for robust methods to address missing data.
@@
Figure 3. Mel-frequency cepstrum and Generalized Cross-Correlation of clean and corrupted audio.
Figure 4. Effect of MDP on the $E_{DoA}$
Figure 5. Effect of trivial methods on the $E_{DoA}$
Challenges in Replacing Corrupted Channels. Given the high correlation between closely spaced microphones, one might assume that excluding a faulty channel is straightforward. However, experiments show that this assumption is not effective in practice.
We conducted an experiment with two scenarios with a 75% MDP. First, we replaced missing data in faulty channels with information from the most correlated channel ($REP_{corr}$). In the second scenario, we excluded faulty channels altogether ($EXC$) and retrained the DNN to work only with the remaining active channels.
Figure  5 shows that both methods increased DoA error, by 46.34% and 63.02%. These findings demonstrate that neither detection nor retraining with fewer channels is sufficient to handle the dynamic nature of intermittent power systems, where missing data varies unpredictably. Traditional approaches fail to manage this challenge effectively in real-world applications. A more comprehensive, information entropy-based approach is needed to address missing data effectively, without requiring prior knowledge of corrupted microphones or relying on time-domain recovery models.
@@ 3. Problem Formulation and Mathematical Foundation
This section formulates the problem and provides the mathematical foundation of the proposed LOCUS.
First, we revisit the definitions of Full Rank and Low Rank Matrices. A full rank matrix $F$ has all its rows and columns linearly independent. In the context of feature extraction for sound source localization, this means that all channels provide complete and accurate information, ensuring no data loss. Conversely, a low-rank matrix $\tilde{F}$ has rows or columns that are linearly dependent, indicating missing or corrupted information in one or more channels. This loss or corruption of data reduces the effectiveness of the sound source localization algorithm.
Consider $X^{n\times T}$ as the raw time-domain data representing the input multichannel audio from a microphone array. $X^{n\times T}$ is a set of $n$ random variables , $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, where $n$ is the number of input channels and $T$ is the time duration of each channel. A feature extraction function, $G$, maps the time-domain data into the feature domain, $F^{n\times D}=G(X)$, where $D$ is the length of each feature.
When all information from all channels is present, $F$ is considered a full rank matrix. However, when one or more input channels have missing or corrupted elements, $X^{n\times T}$ changes to $\tilde{X}^{n^{\prime}\times T}$, where $n^{\prime}$ is less than $n$. As a result, the feature extraction function generates $\tilde{F}^{n^{\prime}\times D}=G(\tilde{X})$. In this paper, we focus on recovering $\hat{F}$ from $\tilde{F}$, where $\hat{F}\approx F$.
To recover the missing information, we use interpolation based on the existing information in $\tilde{F}$. However, during interpolation, the existing information in the interpolated feature map, $\bar{\bar{F}}$, may get polluted, introducing additional errors. To eliminate this residual error, we use the less-affected input information as feedback to the estimated feature.
Since it is unknown which portions of $\tilde{F}$ contain available information, we first identify the information content of $\tilde{F}$. Given that $\tilde{F}$ is a function of $\tilde{X}$, the information entropy at the $(j,k)^{\text{th}}$ element of $\tilde{F}$ is formulated as:
@@
Here, the sigmoid function $\sigma$ in Equation 1 normalizes the entropy value between 0 and 1. $I_{j,k}(\tilde{F})$ is a real number between 0 and 1, where 1 represents the presence of all information in the $(j,k)^{\text{th}}$ element of $\tilde{F}$, while 0 indicates a complete absence of information in that element. Any value between 0 and 1 implies fractional missing information. For a full-rank feature map $F$, every element of $I(F)$ is 1.
The recovered feature set, $\hat{F}$, can be written as:
@@
Here, $\otimes$ represents element-wise multiplication, and $I(\tilde{F})\otimes(\bar{\bar{F}}-\tilde{F})$ represents the feedback from the available information of $\tilde{F}$. We can rewrite Equation 2 as:
@@
The first term represents the information content of $\tilde{F}$ written in terms of $\tilde{F}$ and the element-wise information entropy $I(\tilde{F})$. The second term represents the interpolated missing information extracted from $\bar{\bar{F}}$. This paper aims to determine the estimated full-rank feature matrix, $\hat{F}$, from a given low-rank feature matrix $\tilde{F}$.
4. LOCUS: Entropy-Guided Feature Restoration Framework
Figure 6. Architecture of the proposed LOCUS. The highlighted components are the contributions of this paper.
Our proposed LOCUS recovers the perturbed information of the low-rank feature matrix, $\tilde{F}$ (Equation 3), and estimates the full-rank feature matrix, $\hat{F}\approx F$. Figure 6 shows the architecture of LOCUS, which sits between the feature extraction module and the DNN-based sound source localization (SSL) task. LOCUS has three main components: (1) Early Attention (InFo), which estimates the element-wise information entropy, $I(\tilde{F})$, indicating the amount of available information in $\tilde{F}$; (2) Deep Conditional Interpolator (LaFS), which interpolates the feature set, $\bar{\bar{F}}$, from the available low-rank feature matrix, $\tilde{F}$; and (3) Guided Replacement (GRep), which combines the available information from the original feature set with the interpolated feature values guided by element-wise information entropy. Equation 3 governs this combining process. The resultant estimated full-rank feature set, $\hat{F}$, is then fed to the downstream SSL task for inference.
4.1. Information-Weighted Focus (InFo)
The purpose of Information-Weighted Focus (InFo) module is to estimate the information entropy, $I(\tilde{F})$, for each element of the low-rank feature set $\tilde{F}$. The InFo module estimates the channel and spatial information entropy (Park et al., 2018) by giving higher attention to elements with more available information. The input to InFo is the low-rank feature set $\tilde{F}\in\mathbb{R}^{d_{\tilde{X}}}$, where $d_{\tilde{X}}$ is the dimension of $\tilde{F}$. The left-hand side of Figure 6 shows the proposed early attention architecture, consisting of two main components.

### v3

@@ Abstract.
Accurate sound source localization (SSL), such as direction-of-arrival (DoA) estimation, relies on consistent multichannel data. However, batteryless systems often suffer from missing data due to the stochastic nature of energy harvesting, degrading localization performance. We propose LOCUS, a deep learning framework that recovers corrupted features in such settings. LOCUS integrates three modules: (1) Information-Weighted Focus (InFo) to identify corrupted regions, (2) Latent Feature Synthesizer (LaFS) to reconstruct missing features, and (3) Guided Replacement (GRep) to restore data without altering valid inputs. LOCUS significantly improves DoA accuracy under missing-channel conditions, achieving up to 36.91% error reduction on DCASE and LargeSet, and 25.87–59.46% gains in real-world deployments. We release a 50-hour multichannel dataset to support future research on localization under energy constraints. Our code and data are available at: {https://bashlab.github.io/locus_project/}
Sound Source Localization (SSL), Feature recovery, Batteryless embedded systems, Deep Neural Network (DNN).
ccs: Computer systems organization Sensors and actuatorsccs: Computer systems organization Embedded software
@@ 1. Introduction
Direction-of-arrival (DoA) estimation enables sound source localization (SSL) by extracting spatial and temporal cues from multichannel microphone data. Classical signal processing methods (Schmidt, 1986; DiBiase, 2000) and deep learning approaches (Yalta et al., 2017; Chakrabarty and Habets, 2017) have been applied to DoA estimation, with deep learning offering better robustness to real-world challenges like multipath and reverberation.
Intermittently powered batteryless systems face challenges in acquiring full-channel data due to the stochastic nature of energy harvesting. These systems operate using capacitors that store harvested energy, enabling brief active phases followed by sleep when energy depletes. While capacitor size can control these cycles, longer events often span multiple sleep phases, causing data loss (Monjur et al., 2023). Since DoA estimation requires consistent multichannel input, these interruptions severely degrade localization accuracy. As a result, batteryless systems are typically limited to short classification tasks (Majid et al., 2020a; Islam and Nirjon, 2020; Monjur et al., 2023), and rarely applied to DoA estimation. Still, their low-power operation holds strong potential for applications like environmental monitoring and wildlife tracking, where source localization is essential.
Existing methods for recovering missing data in audio streams rely on statistical imputation—mean (Allison, 2001), hot deck (Little and Rubin, 2019), or multiple imputations (Rubin, 1996)—but suffer from oversimplification, mismatch sensitivity, or high computational cost (Hasan et al., 2021). Mutual information-based techniques (Vergara and Estévez, 2014) can detect corrupted regions but fail to recover features effectively for DNN-based DoA estimation.
Recent DNN-based imputation pipelines for batteryless systems (Monjur et al., 2023) address data recovery but are limited to single-channel input, leaving multichannel recovery for DoA estimation an open challenge. Instead of reconstructing raw audio, prior works suggest that recovering key feature representations—such as generalized cross-correlation (GCC) and Mel– Frequency Cepstral Coefficients (MFCC) is more power-efficient and preserves spatial localization cues (de Godoy et al., 2018; Gobieski et al., 2018). This approach is particularly well-suited for batteryless devices, where transmission cost dominates power consumption.
Batteryless devices offer a compelling solution for long-term, maintenance-free sensing in challenging or resource-constrained settings. Motivated by the need for spatial awareness in such scenarios, we design our system to perform DoA estimation under the energy and data constraints typical of batteryless platforms, making spatial intelligence feasible where it was previously impractical. To enable robust DoA estimation in batteryless deployments with intermittent multichannel data loss, we propose LOCUS, an offline deep learning framework that reconstructs corrupted multichannel features for sound source localization. Rather than recovering raw audio—which is power-intensive to transmit and process—LOCUS focuses on reconstructing key feature representations.
LOCUS comprises three modules: (1) Information-Weighted Focus (InFo), which estimates entropy to locate unreliable features; (2) Latent Feature Synthesizer (LaFS), which reconstructs missing values based on learned relationships; and (3) Guided Replacement (GRep), which replaces only corrupted elements while preserving valid data. The reconstructed features are then passed to a downstream DNN for localization. While the components used in LOCUS are studied separately in previous literature, their integration under the same modular framework enables existing SSL models to function robustly under non-ideal conditions, without requiring changes to SSL model architecture. Our contributions could be summarized as follows.
  • We propose LOCUS, a three-stage framework that recovers corrupted multichannel features using entropy-guided interpolation, enabling robust localization in batteryless systems.
  • We evaluate LOCUS across audio and non-audio tasks—on benchmark datasets, and real-world solar and RF-powered environments—and release LargeSet, a 50-hour multichannel dataset for future research.
While many approaches address sparse data recovery, LOCUS is unique in targeting feature-domain recovery under real-world spatial and temporal dropout in multichannel arrays. Unlike compressed sensing or matrix completion methods, LOCUS supports partial observability, asynchronous channels, and energy-constrained operation. Its physically inspired attention modules capture spatial cues, which are essential for sound localization capabilities and are not modeled by general-purpose recovery methods.
LOCUS achieves up to 21.27% and 17.14% lower DoA error than $SoundSeive$ (Monjur et al., 2023) on DCASE (Politis et al., 2021) and LargeSet, respectively, and improves real-world performance by 25.87–59.46% under stochastic dropout. It also reduces 90th percentile localization error by 25.52% in Wi-Fi CSI-based distance estimation (Ayyalasomayajula et al., 2020). LOCUS is well-suited for low-power deployments such as smart buildings and wildlife monitoring, where spatial feature recovery—rather than waveform reconstruction—is sufficient for accurate localization.
2. Missing Data in Batteryless Microphone Arrays
Missing Channels in Intermittent Systems. Batteryless systems powered by RF, solar, or vibration harvesters often experience unpredictable power fluctuations, leading to data loss in multi-channel microphone arrays. Environmental dynamics—like human movement, light occlusion, or mechanical interference—cause inconsistent power, resulting in intermittent microphone failures. In distributed setups, spatial variations in harvested energy lead to asynchronous channel availability, while monolithic systems may deactivate subsets of microphones to conserve power. Although full dropout is possible, partial and asynchronous channel loss is far more common and presents a recurring challenge for localization. As shown in Section 7, these power variations significantly degrade localization accuracy. LOCUS addresses this by reconstructing features and estimating source locations offline, making it ideal for batteryless and low-power systems without requiring on-device inference.
Quantifying Missing Data (MDP). We define a channel as the full signal from a single microphone in a microphone array. A channel is considered missing if its entire data stream is absent due to energy loss or simulated dropout. To quantify missingness, we define the Missing Data Percentage (MDP) as a temporal metric, where the fraction of time during which one or more channels are unavailable:
@@ 2. Missing Data in Batteryless Microphone Arrays
where $t_{1},t_{2},\ldots,t_{n}$ are the time segments with one or more missing channels, and $T$ is the total audio duration.
Figure 1. Example of raw audio from a four-channel microphone array. Each row represents one microphone; gray regions indicate missing data (75% MDP).
Figure 1 shows an example with 75% MDP: during segments $t_{1}$ and $t_{2}$, at least one channel is missing, while $t_{3}$ has full data across all microphones. Since 75% of the total duration includes missing channels, the MDP is 75%. The temporal MDP is preferred over a channel-count-based MDP because it aligns with realistic, intermittent dropout patterns encountered in deployment, capturing both timing and duration of corruption, and enabling consistent evaluation across synthetic and real scenarios.
@@
Figure 2. Comparison of multi-channel features with and without missing data. Left: Clean MFCC and GCC; Right: Corrupted versions with missing channels.
Figure 3. DoA error ($E_{DoA}$) increases as the Missing Data Percentage (MDP) rises. Results shown for a baseline SSL model evaluated on two datasets (details in Section 5.1) under varying levels of missing data.
Impact of Missing Channels on SSL. While SSL has been widely studied in both simulated and real-world settings (Evers et al., 2020; Tan et al., 2021), the impact of missing channel data remains underexplored. Key features like Mel-frequency cepstral coefficients (MFCC) and Generalized Cross-Correlation (GCC) are highly sensitive to missing inputs, distorting the feature space and degrading DNN performance. As shown in Figure 2, missing channels significantly alter MFCC and GCC patterns. We quantify this by evaluating a standard SSL model (Shimada et al., 2021; Adavanne et al., 2018) on two datasets under increasing MDP levels; DoA error rises sharply (Figure 3), underscoring the need for more robust recovery methods.
Figure 4. DoA error increases when using naive recovery methods: replacing with the most correlated channel ($REP_{corr}$) or excluding faulty channels ($EXC$).
Challenges in Replacing Corrupted Channels. Although nearby microphones are often correlated, naive approaches such as channel exclusion or correlated-channel replacement fail under dynamic dropout. We tested two such methods at 75% MDP: $REP_{corr}$, which replaces missing data with the most correlated working microphone’s signal; and $EXC$, which excludes missing channels and retrains the DNN on only the active ones. As shown in Figure 4, both approaches increased DoA error by 46–63%. These results highlight the need for adaptive strategies like entropy-based recovery, which operate without prior knowledge of which channels are corrupted.
@@ 3. Problem Formulation and Mathematical Foundation
We begin by revisiting the concepts of Full-Rank and Low-Rank Matrices. A full-rank matrix $F$ consists of linearly independent rows and columns, implying that all channels contribute distinct and complete information—critical for accurate sound source localization. In contrast, a low-rank matrix $\tilde{F}$ contains linearly dependent rows or columns, typically arising from missing or corrupted channel data, thereby degrading localization performance.
Let $X^{n\times T}$ represent the raw multichannel time-domain signal, where $n$ denotes the number of channels and $T$ is the time duration per channel. We define $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ as the set of input signals. A feature extraction function $G$ maps this input to a feature domain as $F^{n\times D}=G(X)$, where $D$ is the feature length per channel.
When all channels are present and functioning, $F$ is full-rank. If one or more channels are missing or degraded, we denote the corrupted input as $\tilde{X}^{n^{\prime}\times T}$ with $n^{\prime}<n$, and its corresponding features as $\tilde{F}^{n^{\prime}\times D}=G(\tilde{X})$. Our objective is to recover an estimate $\hat{F}$ that closely approximates the full-rank feature matrix $F$, i.e., $\hat{F}\approx F$.
To achieve this, we initially interpolate the missing features from $\tilde{F}$, yielding $\bar{\bar{F}}$. However, this interpolation can inadvertently distort the original information, introducing additional error. To mitigate this, we reintroduce the relatively more reliable parts of $\tilde{F}$ as feedback into the interpolated features.
Because the reliability of individual elements in $\tilde{F}$ is unknown, we introduce a heuristic confidence mask $I(\tilde{F})$, which estimates the presence or absence of information. This mask is computed for each element $(j,k)$ of $\tilde{F}$ as:
@@
Here, $\sigma$ is the sigmoid function that bounds the confidence values between 0 and 1. This formulation is not derived from a formal probabilistic model, but rather serves as a heuristic measure of reliability. A value of 1 indicates high confidence in the information content of $\tilde{F}_{j,k}$, while 0 indicates complete absence, with intermediate values suggesting partial reliability.
Using this confidence mask, we refine the interpolated feature map as:
@@
This expression can be equivalently rewritten as:
@@
In this formulation, $I(\tilde{F})$ acts as a confidence-driven selector: it retains elements from $\tilde{F}$ where the confidence is high and replaces uncertain elements with interpolated values from $\bar{\bar{F}}$. Our goal throughout this paper is to estimate a full-rank feature matrix $\hat{F}$ from a corrupted, low-rank feature matrix $\tilde{F}$ using this reliability-guided reconstruction process.
4. LOCUS: LOcalization with Channel Uncertainty and Sporadic Energy
Figure 5. LOCUS consists of three key components for robust feature recovery under missing data. InFo estimates element-wise confidence score, $I(\tilde{F})$, from the low-rank feature map $\tilde{F}$ using both channel-wise and spatial feature branches. LaFS reconstructs a full-rank feature representation $\bar{\bar{F}}$ by modeling dependencies across available channels via a symmetric encoder-decoder structure. GRep combines the original and synthesized features using $I(\tilde{F})$ to produce the final recovered feature map $\hat{F}$.
LOCUS estimates the full-rank feature matrix $\hat{F}\approx F$ from a corrupted low-rank matrix $\tilde{F}$, following the recovery formulation in Equation 3. As shown in Figure 5, LOCUS operates between the feature extractor and the DNN-based sound source localization (SSL) model. It consists of three key components: (1) Information-Weighted Focus (InFo), which estimates element-wise confidence score $I(\tilde{F})$; (2) Latent Feature Synthesizer (LaFS), which reconstructs a complete feature set $\bar{\bar{F}}$ from $\tilde{F}$; and (3) Guided Replacement (GRep), which fuses $\tilde{F}$ and $\bar{\bar{F}}$ using confidence masking $I(\tilde{F})$ to produce $\hat{F}$. This process is governed by Equation 3, and the resulting $\hat{F}$ is passed to the downstream SSL model for inference.
4.1. InFo: Information-Weighted Focus
The purpose of Information-Weighted Focus (InFo) module is to estimate the confidence score, $I(\tilde{F})$, for each element of the low-rank feature set $\tilde{F}$. The InFo module estimates the channel and spatial confidence score  (Park et al., 2018) by giving higher attention to elements with more available information. The input to InFo is the low-rank feature set $\tilde{F}\in\mathbb{R}^{d_{\tilde{X}}}$, where $d_{\tilde{X}}$ is the dimension of $\tilde{F}$. The left-hand side of Figure 5 shows the proposed early attention architecture, consisting of two main components.